OpenStudy (anonymous):

Solve: w^2-7w-8=0 it ask what the solution is for w

OpenStudy (anonymous):

(w-8)(w+1)=0 w=8 w=-1

OpenStudy (anonymous):

Does it matter how it is written (-1,8) or (8,-1)

OpenStudy (anonymous):

(-b(+-)sqrt(b^2-4ac))/2a. a=w^2 = 1 b=-7w = -7 c=-8 w= (7+-sqrt(49-4*(1*-8)))/2 = (7+-sqrt(81))/2 = (7+-9)/2 = 16/2 and -2/2 => w=8 V w=-1 It doesn't matter which order you write the answers.

OpenStudy (anonymous):

V means "and" in mathematical notation. Just wanted to avoid any confusion.

OpenStudy (anonymous):

oh ok